The  domain of the function f / g where f (x, y) = 2xy and

g(x, y) = x ²+ y² is R² .is it true or false. Give reasons for your answer


1
Expert's answer
2021-02-03T01:17:14-0500

First of all, as "f,g" are polynomial, "\frac{f}{g}" are well defined on "\mathbb{R}^2 \setminus \{g=0\}". The only point where "g(x,y)=0" is "x=y=0". Thus, the function is not defined on "(0;0)". In addition, we can not extend "f\/g" to (0,0) in any natural way, as the limit "\lim_{x,y\to 0} \frac{f(x,y)}{g(x.y)} = \lim_{r\to 0} \frac{2 r^2 \sin \theta \cos \theta}{r^2}" does not exist (as it depends on "\theta"). Therefore "\frac{f}{g}" is not defined on "\mathbb{R}^2" and can't be extended in any natural way.